ok I have major homework and this math is really hard so will u please help me,
Steven is 14 years old. His mom is 38. How many years ago was the mom exactly 5 times as old as Steve?
number of years?
Steven then?
mom then?

Answer by Paul(988)

Steven's age =14years
His mom's age = 38
Let the age be x
To determine the age of Steven so x-14 also this would affect his mom's age so x-38
Equation:
5(x-14)=x-38
5x-70=x-38
4x=32
x=8

Hence, 8 years ago his mom was 5 times as old as him